# Univ. Montpellier, IMT Mines Alès, Montpellier, France

### v1

@@ 1 Introduction
The remainder of this paper is structured as follows. Section 2 presents the theoretical foundations, encompassing belief functions, evidential clustering, and explainability. Section 3 introduces utility functions, representativeness, evidential mistakeness, and the IEMM algorithm. We provide formal analysis of these concepts along with illustrative examples. We also make available the complete code for all experiments at (OMITTED TO AVOID IDENTIFICATION).
@@ 2.1 Belief Functions
  • A mass function with exactly one focal element $A$ is called categorical, representing the logical assertion that ’$\omega$ belongs to $A$’. If this single focal element is $\Omega$ itself, the function is vacuous, conveying no information beyond the closed-world hypothesis.
@@ 3.1 Explaining Classifiers under Uncertainty and Imprecision
Let us consider a categorical evidential classifier [^2] We use the term classifier to emphasize that the development of this section is valid not only for clustering but also for all (categorical) evidential partitions of the data. $\mathcal{M}_{c}:X\rightarrow\mathbb{M}$. We define ¯$\overline{\mathcal{M}}_{c}:X\rightarrow 2^{\Omega}$ such that, for any $x\in X$ and $A\subseteq\Omega$, $\mathcal{M}_{c}(x)(A)=1$ if and only if ¯$\overline{\mathcal{M}}_{c}(x)=A$.
@@
While in the hard case the DT explainer quality relates closely to representativity (details in appendix 0.B), in the evidential setting this concept needs refinement since errors vary in severity. For instance, assigning an observation $x$ with ¯$\overline{\mathcal{M}}_{c}(x)=\{\omega_{1}\}$ to $\{\omega_{1},\omega_{2}\}$ is potentially less severe than assigning it to $\{\omega_{2}\}$ alone. To capture these nuances, we introduce the concept of utility that quantifies the satisfaction of a decision-maker when assigning a subset $A$ when the ground truth is $B$.
@@
$\underline{\operatorname{Cost}_{\mathcal{M}_{c},\Delta}}(x)$ 00 $1-\mathcal{U}(\{\omega_{1},\omega_{2}\},\{\omega_{1}\})$ $1$
@@
$\overline{\operatorname{Cost}_{\mathcal{M}_{c},\Delta}}(x)$ $\mathcal{U}(\{\omega_{1},\omega_{2}\},\{\omega_{1}\})$ $1$ $1+\mathcal{U}(\{\omega_{1},\omega_{2}\},\{\omega_{1}\})$
Table 1: For each point highlighted in Figure 4, we present the point $x$, the metacluster assigned by classifier $\Delta$, the cost of assigning $x$ to the metacluster designated by $\Delta$, the set of metaclusters not assigned by $\Delta$, and the cost of not assigning $x$ to them. Note that ¯$\overline{\mathcal{M}}_{c}(x_{0})=\overline{\mathcal{M}}_{c}(x_{1})=\overline{\mathcal{M}}_{c}(x_{2})=\{\omega_{1}\}$.
@@
$\overline{M}_{\mathcal{M},\mathcal{U}}(S)=\sum_{x\in S}\overline{\operatorname{Cost}_{\mathcal{M},\Delta}}(x)\text{ and }\underline{M}_{\mathcal{M},\mathcal{U}}(S)=\sum_{x\in S}\underline{\operatorname{Cost}_{\mathcal{M},\Delta}}(x).$
@@ 3.2 Implementation
When an explainer yields a metacluster $A$, while the original classifier assigns $B$, two types of errors can occur. The first is insufficient coverage, measured by $|A^{C}\cap B|$ where the explainer fails to include all elements of the true metacluster. The second is excessive coverage, measured by $|A\cap B^{C}|$ where the explainer includes elements not in the true metacluster. Penalizing insufficient coverage indicates the explainer is not cautious enough, while penalizing excessive coverage suggests it is too cautious.
@@
$\overline{\mathcal{U}}^{\lambda}(A,B)=\left(\frac{|A\cap B|}{|A\cup B|}\mathbbm{1}_{B\subset A}\right)^{1/\lambda}\text{ and }\underline{\mathcal{U}}^{\lambda}(A,B)=\left(\frac{|A\cap B|}{|A\cup B|}\mathbbm{1}_{A\subset B}\right)^{1/\lambda}.$
@@
[…]
@@
1:$\Delta\leftarrow\textbf{split\_tree}(X,\mathcal{M}(X),\mathbb{F},v)$
@@
16:     node.condition "$x_{i}\leq\theta"$
17:     node.lt $\leftarrow$ split_tree($(\{x^{j}\}_{j\in L},\{m^{j}\}_{j\in L},F_{L},v)$
18:     node.rt $\leftarrow$ split_tree($(\{x^{j}\}_{j\in R},\{m^{j}\}_{j\in R},F_{R},v)$
@@
We have implemented IEMM using Python 3.9.6. All code is available at (OMITTED TO AVOID IDENTIFICATION). The implementation of a decision tree accepting evidential labels was based on the code made available by [17].
@@ Appendix 0.A Appendix: On Simplification Explanation Techniques
[…]

### v2

@@ 1 Introduction
The remainder of this paper is structured as follows. Section 2 presents the theoretical foundations, encompassing belief functions, evidential clustering, and explainability. Section 3 introduces utility functions, representativeness, evidential mistakeness, and the IEMM algorithm. We provide formal analysis of these concepts along with illustrative examples. We also make available the complete code for all experiments at https://github.com/victorsouza89/iemm.
@@ 2.1 Belief Functions
  • A mass function with exactly one focal element $A$ is called categorical, representing the logical assertion that $\omega$ belongs to $A$’. If this single focal element is $\Omega$ itself, the function is vacuous, conveying no information beyond the closed-world hypothesis.
@@ 3.1 Explaining Classifiers under Uncertainty and Imprecision
Let us consider a categorical evidential classifier [^2] We use the term classifier to emphasize that the development of this section is valid not only for clustering but also for all (categorical) evidential partitions of the data. $\mathcal{M}_{c}:X\rightarrow\mathbb{M}$. We define $\overline{\mathcal{M}}_{c}:X\rightarrow 2^{\Omega}$ such that, for any $x\in X$ and $A\subseteq\Omega$, $\mathcal{M}_{c}(x)(A)=1$ if and only if $\overline{\mathcal{M}}_{c}(x)=A$.
@@
While in the hard case the DT explainer quality relates closely to representativity (details in appendix 0.B), in the evidential setting this concept needs refinement since errors vary in severity. For instance, assigning an observation $x$ with $\overline{\mathcal{M}}_{c}(x)=\{\omega_{1}\}$ to $\{\omega_{1},\omega_{2}\}$ is potentially less severe than assigning it to $\{\omega_{2}\}$ alone. To capture these nuances, we introduce the concept of utility that quantifies the satisfaction of a decision-maker when assigning a subset $A$ when the ground truth is $B$.
@@
$\underline{\operatorname{Cost}_{\mathcal{M}_{c},\Delta}}(x)$ 0 $\mathcal{U}(\{\omega_{1},\omega_{2}\},\{\omega_{1}\}) 11$
@@
$\overline{\operatorname{Cost}_{\mathcal{M}_{c},\Delta}}(x)$ $\mathcal{U}(\{\omega_{1},\omega_{2}\},\{\omega_{1}\})$ 11 $1+\mathcal{U}(\{\omega_{1},\omega_{2}\},\{\omega_{1}\})$
Table 1: For each point highlighted in Figure 4, we present the point $x$, the metacluster assigned by classifier $\Delta$, the cost of assigning $x$ to the metacluster designated by $\Delta$, the set of metaclusters not assigned by $\Delta$, and the cost of not assigning $x$ to them. Note that $\overline{\mathcal{M}}_{c}(x_{0})=\overline{\mathcal{M}}_{c}(x_{1})=\overline{\mathcal{M}}_{c}(x_{2})=\{\omega_{1}\}$.
@@
[…]
@@ 3.2 Implementation
When an explainer yields a metacluster $A$, while the original classifier assigns $B$, two types of errors can occur. The first is insufficient coverage, measured by $|A^{C}\cap B|$ - where the explainer fails to include all elements of the true metacluster. The second is excessive coverage, measured by $|A\cap B^{C}|$ - where the explainer includes elements not in the true metacluster. Penalizing insufficient coverage indicates the explainer is not cautious enough, while penalizing excessive coverage suggests it is too cautious.
@@
[…]
@@
We also define special cases as limits when $\lambda$ approaches 0 and $\pm\infty$. That gives $\mathcal{U}^{0}(A,B)=\mathbbm{1}_{A=B},\ \mathcal{U}^{-\infty}(A,B)=\mathbbm{1}_{A\subset B}\text{ and }\mathcal{U}^{\infty}(A,B)=\mathbbm{1}_{B\subset A}.$
@@
[…]
@@
16:  node.condition $\leftarrow"x_{i}\leq\theta$"
[…]
@@
We have implemented IEMM using Python 3.9.6. All code is available at https://github.com/victorsouza89/iemm. The implementation of a decision tree accepting evidential labels was based on the code made available by [17].
@@ Appendix 0.A Appendix: On Simplification Explanation Techniques
A Concrete Example: Consider a philosopher studying living beings who observes two key characteristics: their appearance and their mode of locomotion. To formalize this classification problem, the philosopher defines the feature space of conceivable living beings as $\mathbb{X}=\texttt{App}\times\texttt{Move},$ where $\texttt{App}=\{\texttt{feathered},\texttt{featherless}\}\text{ and }\texttt{Move}=\{\texttt{biped},\texttt{non-biped}\}.$ From these features, we can construct the set of feature literals: $\mathbf{L}=\{\langle\texttt{App},\texttt{feathered}\rangle,\langle\texttt{App},\texttt{featherless}\rangle,\langle\texttt{Move},\texttt{biped}\rangle,\langle\texttt{Move},\texttt{non-biped}\rangle\}.$ The set of all consistent subsets of feature literals encompasses all possible combinations that do not contain contradictory values for the same attribute: $\mathbf{C}=\{\varnothing,$ $\{\langle\texttt{App},\texttt{feathered}\rangle\},\{\langle\texttt{App},\texttt{featherless}\rangle\},\{\langle\texttt{Move},\texttt{biped}\rangle\},\{\langle\texttt{Move},\texttt{non-biped}\rangle\},$ $\{\langle\texttt{App},\texttt{feathered}\rangle,\langle\texttt{Move},\texttt{biped}\rangle\},\{\langle\texttt{App},\texttt{feathered}\rangle,\langle\texttt{Move},\texttt{non-biped}\rangle\},$ $\{\langle\texttt{App},\texttt{featherless}\rangle,\langle\texttt{Move},\texttt{biped}\rangle\},\{\langle\texttt{App},\texttt{featherless}\rangle,\langle\texttt{Move},\texttt{non-biped}\rangle\}$ $\}.$ An example of a DNF is given by $\mathbf{D}=\{\{\langle\texttt{App},\texttt{featherless}\rangle,\langle\texttt{Move},\texttt{biped}\rangle\}\}$, which corresponds to featherless bipedal beings. That is, for any living being $x$, we have $\operatorname{DNF}_{\mathbf{D}}(x)=(x_{\texttt{App}}=\texttt{featherless})\land(x_{\texttt{Move}}=\texttt{biped})$. $\operatorname{DNF}_{\mathbf{D}}(x)$ is true whenever $x$ is a human being. Conversely, $\mathbf{D}^{\prime}=\{\{\langle\texttt{App},\texttt{feathered}\rangle\},\{\langle\texttt{Move},\texttt{non-biped}\rangle\}\}$ corresponds to beings that are either feathered or non-bipedal. In this case, the induced DNF is $\operatorname{DNF}_{\mathbf{D}^{\prime}}(x)=(x_{\texttt{App}}=\texttt{feathered})\lor(x_{\texttt{Move}}=\texttt{non-biped})$ and $\operatorname{DNF}_{\mathbf{D}^{\prime}}(x)$ is false whenever $x$ is a human being.